1. Consider a ball of mass 0.050 kg release from 0.85 m height, what velocity does the ball will attain when it hits the floor?


2. In a bouncing ball lab experiment what do you expect about the bounce height of the

ball?


3. Without air resistance, the ball is still not be able to bounce to its original drop height,

why?


4. Name some factors that would affect the bounce height.


1
Expert's answer
2021-06-29T02:12:04-0400

1)


"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(0.85)}=4.0\frac{m}{s}"

2) As the ball goes back up, kinetic energy (now a bit less) gets traded back for gravitational potential energy, and it will rise back to a height that is the original height times (1-fraction of energy lost).


3) Without other forces, such as air resistance, the ball is still not be able to bounce to its original drop height because of these energy transfers (elastic potential energy, thermal energy, sound energy). The ball will never have as much kinetic energy as it originally had


4) The combination of the material properties of a ball (surface textures, actual materials, amount of air, hardness/ softness, and so on) affects the height of its bounce.